Question 3 Lenberg, Inc. believes in the residual dividend policy. This years earnings are expected to total $10,000,000. Lenberg, a very conservative company, is financed solely with common stock. The required rate of return on retained earnings is 12 percent, whereas the cost of newly raised capital is 14 percent because of issuance costs. 12 If Lenberg has $6,000,000 of investment projects with expected returns greater than 12 percent, compute and explain the total amount of dividends that Lenberg pays. If Lenberg has $12,000,000 of investment projects with expected returns greater than 14 percent, compute and explain the total amount of dividends that Lenberg pays If most of the investors of Lenberg are prohibited from spending capital, what kind of dividend policy is most suitable Lenberg? Explain your answer in detail. a. 14 b. pension funds and endowment trusts which are for c. Question4 Gary deposits $100,000 into his futures trading account in brokerage house. He wants to take a short position in Hang Seng Index futures (HSIF) to hedge his securities portfolio risk. Currently, the spot month HSIF is now trading at 25,740. The initial margin requirement is $84,500 and maintenance margin is $67,600. The contract multiplier is $50 per index point. a. How many futures contracts can Gary trade if his account has a cash balance of $100,000? Compute the price of HSIF so that Gary will receive margin call. b.
Answer #1

Question 3

Part (a)

Required Return on retained earnings, RRE = 12%

Return on the investment project, RIP > RRE = 12%

Hence, the company should undertake the project and deploy the earnings to fund this project.

Net Income = 10,000,000

Investment required = 6,000,000

Target capital structure for financing = 100% with equity (common stock)

Hence, net income deployed towards financing the new project = min (Net income, investment required x proportion of equity) = min (10000000, 6000000 x 100%) = 6,000,000

Hence, residual income that can be distributed as dividend = Net income - income deployed towards investment project = 10,000,000 - 6,000,000 = $ 4,000,000

Part (b)

Required Return on retained earnings, RRE = 12%

Required Return on fresh equity capital to be raised, RFE = 14%

Return on the investment project, RIP > 14%, thus RIP > RRE = 12% and RIP > RFE = 14%

Hence, the company should undertake the project and deploy the earnings to fund this project. If there is shortfall, it should fund the project from external equity capital.

Net Income = 10,000,000

Investment required = 12,000,000

Target capital structure for financing = 100% with equity (common stock)

Hence, net income deployed towards financing the new project = min (Net income, investment required x proportion of equity) = min (10000000, 12000000 x 100%) = 10,000,000

Hence, residual income that can be distributed as dividend = Net income - income deployed towards investment project = 10,000,000 - 10,000,000 = 0

Part (c)

Since the shareholders of the Lenberg are prohibited from spending capital, they will ideally want higher dividend yield and high dividend income from the investee companies (like Lenberg) so that they can use the same towards capital expenditures. Hence, Lenberg should then be required to have a liberal dividend policy of paying out say, a constant ratio of net income every year as dividend.